40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 24 04:13:15 2013, Jenne, Update, LSC, PRMI + 2 ALS arms 
    Reply  Thu Oct 24 08:04:54 2013, Steve, Update, LSC, PRMI + 2 ALS arms 
    Reply  Thu Oct 24 09:08:27 2013, jamie., Update, LSC, PRMI + 2 ALS arms 
    Reply  Thu Oct 24 14:14:18 2013, rana, Update, LSC, PRMI + 2 ALS arms 
       Reply  Thu Oct 24 15:19:58 2013, Koji, Update, LSC, PRMI + 2 ALS arms 
    Reply  Thu Oct 24 23:25:37 2013, Jenne, Update, LSC, End transmission triggering 
       Reply  Tue Nov 5 16:47:09 2013, Jenne, Update, LSC, End transmission QPDs 
    Reply  Fri Oct 25 04:03:40 2013, Masayuki, Update, LSC, 'scope and spectrum analyser for REFL165 
       Reply  Fri Oct 25 21:36:51 2013, Masayuki, Update, LSC, 'scope and spectrum analyser for REFL165 
          Reply  Mon Oct 28 14:24:01 2013, Masayuki, Update, LSC, 'scope and spectrum analyser for REFL165 
             Reply  Mon Oct 28 18:57:27 2013, Masayuki, HowTo, LSC, read 'scope and spectrum analyser datas 
Message ID: 9274     Entry time: Thu Oct 24 04:13:15 2013     Reply to this: 9275   9276   9279   9286   9289
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI + 2 ALS arms 

[Masayuki, Jenne, Rana]

We have, for the past hour and a few minutes, had PRMI + 2 arms locked.  Yup, that's right, we did it! (We never gave control of the arms to the IR LSC system, so it's kind of cheating, but it was still cool.)

A little after midnight, we felt that the Yarm was behaving well enough that we could give PRMI + 2 arms a try.  So we did.  Probably around 1am-ish, or maybe a little bit before, we had the system locked. 

How did we do it?

* Locked arms in IR to help find green beatnotes.

* Misalign ETMs, lock and align PRMI.

* Misalign PRM.

* Restore ETMs, find arm resonances, then step away (I did +3 counts, which is 29 kHz).

* Restore PRM, lock PRMI.

* Brought Xarm back close to resonance using ALS (-3 counts). It seems like this may not actually have gotten us back to perfect resonance, but that actually made bringing in the other arm easier.

* Brought Yarm back close to resonance using ALS (-3 counts). 

* Turned on Sensing Matrix notches and oscillators (10,000 counts for MICH, actuating on BS and PRM at 562.01 Hz, 200 counts for PRCL actuating on PRM at 564.01 Hz). 

* Stepped arms back and forth to see how things responded.

Notes:

During this process, particularly during the various arm steps, the PRMI lost lock many times.  However, the ALS system never lost lock for either arm, for an hour and a half or so.  Good work, ALS team!!  The PRMI would reaquire lock (sometimes we'd have to undo whatever arm step we just took, to get farther away from resonance) without any intervention.  It seemed that as we came closer to full arm resonance, we were never able to hold PRMI locked.  This is what is instigating some of our investigations for tomorrow.

Also, Rana reported to me that he turned the c1tst model back off, and opened the door(s?) to the ETMY rack to allow more air flow sometime before midnight, which seems to have reduced the rate of the CPU going over 61 microseconds, as well as reduced the number of times the ETMY suspension glitches.  We definitely need to make some changes so that we're not so close to the edge.  This may have been one of the big things that allowed our success tonight.

The transmission PDs at the ends of the arms are saturating around 50 counts (they have gains of 2e-3 so that they are roughly normalized to 1 being the max power in a single arm).  We need to commission the end transmission QPDs. 

All of the signals looked a little ratty, and we heard lots of noise - Rana suggests that we recommission our CARM servo.

ALS beat info:  [Xarm  40.9 MHz,  -11.4 dB], [Yarm  50.5 MHz,  -17.7 dB]

Things to look at tomorrow:

Data!  I should be able to extract sensing matrix information, even though my sensing matrix software isn't totally ready yet.  I know what the oscillators were doing, and I can look at the PD error signals.  We also save the Offsetter numbers, so I can kind of tell what the PRMI+arms situation was. 

Can we tell by looking at the end laser PZT feedback signals whether we're making our arms longer or shorter?  So that we can tell if we're putting on DARM or CARM offsets.

Spectrum and time series of REFL 165 (our PRMI LSC locking PD) to see if we're saturating while we bring the arms into resonance.  Basically, does anything bad happen, particularly since the PD is not a resonant PD, so there are some 1f signals floating around in addition to the 3f signals.  We want to put in a directional coupler after the PD, before the demod board, and send that signal to a spectrum analyzer and a 'scope.  Hopefully we can use the power of the internet to not need to sit in the IFO room saving data as we move the arms around.  Do we need to put bandpass filters on the PD signal before it goes to the demod board?

Optickle model of 1f vs. 3f signals in the different ports, as the CARM offset is reduced.

Violin notches for the arms - should be put into ALS and LSC models.  It looks like the modes are around 631 Hz, but we should check. 

Hardware for end low gain transmission QPDs.

Software (schmidt triggering) for end transmission QPDs.

Modifying / preparing a matrix in the ALS system so that we can give CARM and DARM offsets conveniently.

ELOG V3.1.3-